OpenStudy (anonymous):

DOES ANYONE KNOW HOW TO GRAPH THIS PLZ!! y= sin(x-pi/4)

OpenStudy (anonymous):

\[y=\sin (x-\pi/4)\]

OpenStudy (anonymous):

> Wolfram|Alpha <

OpenStudy (anonymous):

i know that graphs it for you but i need to know how to graph it by hand, thanks anyway @Blitzkrieg

OpenStudy (anonymous):

Alright. There is a specific form. \[y = a \sin (bx - c) + d\]Amplitude = |a| Period = 2π/b Phase Shift = -/b Vertical Shift = d

OpenStudy (anonymous):

Do you know a sin graphs normally?

OpenStudy (anonymous):

*what a sin graph looks like

OpenStudy (anonymous):

yes a wave

OpenStudy (anonymous):

Alright. Here's what a sin graph looks like normally. Remember that it normally has an amplitude of 1, period of 2π, and goes through the origin.